OpenStudy (anonymous):

rob intial weekly pay was $520. After his first year, this was increased by 4%. At the end of his second year, this salary was incrreased by 2%. what is his weekly pay after his 2nd raise?

OpenStudy (anonymous):

wouldnt it be 520x.40?

OpenStudy (anonymous):

520(0.04)=x x+520= weekly pay weeklypay(.02)=x x+weeklypay=answer

OpenStudy (anonymous):

540.80+10.8160? = weekly pay?

OpenStudy (anonymous):

yes 551.616

OpenStudy (anonymous):

520*1.04=540.80 540.80*1.02=551.616

OpenStudy (anonymous):

ok thanks i guess what messed me up was i tried doing 520(.40)

OpenStudy (anonymous):

.40 is forty percent

OpenStudy (anonymous):

.04 is four percent

OpenStudy (anonymous):

.045 is 4.5%

OpenStudy (anonymous):

.455 is 45.4 percent

OpenStudy (anonymous):

45.5 percent

OpenStudy (anonymous):

oh i can see how a small mistake can get and far off from correct one ha

OpenStudy (anonymous):

yes